what is 80 and 78 averaged together
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

what is 80 and 78 averaged together
---
Is it obvious the answer is 79?
-----------------
Ave = (80+78)/2 = 158/2 = 79
===================================
Cheers,
Stan H.

Answer by mananth(16949)

80 +78= 158
...
158/2 = 79